OpenStudy (anonymous):

8/p+5 + 3/p add the fractions

OpenStudy (anonymous):

Common Denominator: p(p+5) 8p/(p(p+5)) + ( (p+5)3)/(p(p+5)) 8p + 3p + 15 11p + 15 (11p + 15)/(p^2 + 5p)

OpenStudy (saifoo.khan):

\[\frac{11 p+15} {p (p+5)} \]

OpenStudy (anonymous):

8/p+5 + 3/p. We want a common denominator. 8p/p(p+5) + 3(p+5)/p(p+5) is a solution. Equivalently, we can add like terms. (8p + 3p + 15)/p(p+5) =(11p + 15)/p(p+5) =(11p + 15)/(p^2+5p) Hope this helps. If it does, please choose good answer. http://www.tutorsean.net for free math videos.